# Wickstead’s conjecture on positive projections and non-representable Banach lattice algebras

### 16 Apr 2026 v1

@@ Abstract.
Research supported by an FPI–UAM 2023 contract (funded by Universidad Autónoma de Madrid) and by grants PID2020-116398GB-I00 and CEX2023-001347-S (funded by MCIN/AEI/10.13039/501100011033).
@@ 4. On the representation of Banach lattice algebras
In this section we use 3.1 to settle the representation problem for Banach lattice algebras in the negative. We begin with two preparatory lemmas: the first records that the range of a positive projection is itself a Dedekind complete vector lattice, and the second shows that disjointness from a positive idempotent in $\mathcal{L}_{r}(X)$ descends to disjointness from the identity on its range. These lemmas will let us apply 3.1 on the range of an arbitrary positive idempotent of $\mathcal{L}_{r}(X)$.
@@
As $a$ ranges over $\{a\in X_{+}:a\leq y\}$, the element $w=Ea$ ranges over the order interval $0,y]_{Y}=\{w\in Y_{+}:w\leq y\}$ in $Y$: indeed, $E$ is positive and $Ea\leq Ey=y$, so $Ea\in[0,y]_{Y}$; and conversely, every $w\in[0,y]_{Y}$ is realized by choosing $a=w$. Therefore
@@
The set whose infimum is taken is the same subset of $Y$ in both (4) and (5). By hypothesis, $(E\wedge T)(y)=0$, so 0 is the greatest lower bound of this set in $X$. Since 0 belongs to $Y$ and the order on $Y$ is inherited from $X$, 0 is also a lower bound of this set in $Y$; and any other lower bound in $Y$ is in particular a lower bound in $X$. Therefore 0 is also the greatest lower bound in $Y$, that is, $(\operatorname{id}_{Y}\wedge T|_{Y})(y)=0$.
@@
If $\alpha\not\in\{0\}\cup\{1/n:n\in\mathbb{N}\}$, then $A$ has no faithful representation on any Dedekind complete vector lattice.

### 20 Apr 2026 v2

@@ Abstract.
Research supported by an FPI–UAM 2023 contract (funded by Universidad Autónoma de Madrid) and by grants PID2024-162214NB-I00 and CEX2023-001347-S (funded by MCIN/AEI/10.13039/501100011033).
@@ 4. On the representation of Banach lattice algebras
In this section we use 3.1 to settle the representation problem for Banach lattice algebras in the negative. We begin with two preparatory lemmas: the first records that the range of a positive projection is itself a Dedekind complete vector lattice, and the second shows that disjointness from a positive idempotent in $\mathcal{L}_{r}(X)$ becomes disjointness from the identity on its range. These lemmas will let us apply 3.1 on the range of an arbitrary positive idempotent of $\mathcal{L}_{r}(X)$.
@@
As $a$ ranges over $\{a\in X_{+}:a\leq y\}$, the element $w=Ea$ ranges over the order interval $0,y]_{Y}=\{w\in Y_{+}:w\leq y\}$ in $Y$. Therefore
@@
The set whose infimum is taken is the same subset of $Y$ in both (4) and (5). The only difference is that the first infimum is taken in $X$ while the second is taken in $Y\subseteq X$. Therefore
$(\operatorname{id}_{Y}\wedge T|_{Y})(y)\leq(E\wedge T)(y)=0.$
@@
If $\alpha\not\in\{0\}\cup\{1/n:n\in\mathbb{N}\}$, then $A$ has no faithful representation on $\mathcal{L}_{r}(X)$, for $X$ any Dedekind complete vector lattice.